# Galaxy And Mass Assembly (GAMA): from filaments to voids, how extreme environment affects gas metallicity and SFR in galaxies

### v1

@@ 2.4 Galaxies in filaments, tendrils, and voids
In this work, we used the classification provided by Alpaslan et al. (2014a), who applied the FoF algorithm (Merson et al. 2013) to classify the galaxies in the GAMA survey. This classification divides galaxies into the following types.
Galaxies in filaments: This classification includes all galaxies within an orthogonal distance of 4.12 Mpc from filaments. Therefore, they are situated in dense environments and are susceptible to the influences of their surroundings.
Tendrils galaxies: Included in this category are all galaxies within an orthogonal distance (4.56 Mpc) from tendrils, and simultaneously located beyond a minimum distance of 4.12,Mpc from any filament. Tendrils are slender, thread-like extensions that serve as connecting links between certain filaments. This term was first introduced by Alpaslan et al. (2014b).
Void galaxies: These are galaxies that are more than 4.56 Mpc away from both filament and tendril structures. As a result, they are situated in largely underdense regions called voids, where they remain unaffected by external influences.
Considering these different types of galaxies, the sample of SF galaxies used in this study has been classified into these three types, by matching the GAMA catalogue and the catalogue of filaments (Alpaslan et al. 2014a), as shown in Table 3.
Environment | Number of SF galaxies |
Filaments | 2 068 |
Tendrils | 1 466 |
@@
From the initial 13 689 SF galaxies, 15% correspond to galaxies in filaments, 11% to tendril galaxies, and 1% to void galaxies. The percentage of remaining unclassified galaxies corresponds mostly to field galaxies and will be used in the following subsection to create the control sample.
@@ 2.5 Control samples
We construct such a control sample using field galaxies, selected to serve as a reference population. Field galaxies are identified from the original GAMA spectroscopic catalogue by excluding all galaxies in filaments, tendrils, and voids. In addition, we also removed galaxies that are members of groups or pairs. These exclusions are based on classifications from the GAMA filament catalogue (Alpaslan et al. 2014a), and the GAMA Galaxy Group Catalogue (G3C) (Robotham et al. 2011). Note that we work only with SF galaxies with the same S/N and redshift cuts described previously in this section. By applying these criteria, we obtain a well-defined and uncontaminated sample of field galaxies for our analysis.
@@
Previous studies investigating environmental effects have generated control samples using similar methodologies (Ellison et al. 2008, e.g.,; Scudder et al. 2012; Chen et al. 2020; Garduño et al. 2021; Sotillo-Ramos et al. 2021) using these criteria. Note that field galaxies are used as a control sample because they are not exposed to extreme environments, that is, they are neither located in structures such as filaments or tendrils, nor highly underdense regions such as voids. This makes them suitable for defining control samples, as done in previous studies (e.g., Ellison et al. 2008).
@@
To ensure the validity of our study, these field galaxies must share similar characteristics with those in filaments, tendrils, and voids. We followed previous works and used the stellar mass and colour as our main criteria (Ellison et al. 2008, e.g.,, 2011, 2013), since these two parameters robustly characterise the main properties of a galaxy (Baldry et al. 2004, Peng et al. 2010). Given that our sample of galaxies is relatively local (z0.3), evolutionary effects are negligible (e.g., Lara-López et al. 2009). Therefore, we ensure that the differences in stellar mass and colour between galaxies in filaments, tendrils, and voids and those in the control sample remain minimal, adhering to specific selection criteria.
@@
These selection criteria, based on stellar mass and colour, were established using the Kolmogorov-Smirnov (KS) and Anderson-Darling (AD) tests as references. These tests assess the null hypothesis that two samples originate from the same population without requiring the specification of the distribution function of that population. The KS test is good enough to verify if the created control sample and the sample of galaxies in filaments/tendrils/voids can be assumed to be drawn from the same galaxy population. However, we complement this test with the AD test since it is more sensitive to the tails of the Empirical Cumulative Distribution Function (ECDF), thus ensuring similarity between the samples. The KS and AD tests accept the null hypothesis that both samples can be assumed to be drawn from the same population if the p-value is higher than the significance level = 0.05.
@@
During the creation of the control sample, additional considerations include limiting the number of matched field galaxies per target galaxy. For galaxies in filaments, a maximum of 20 field galaxies per filament galaxy was set. For tendril galaxies, this limit ranges from 2 to 20, while for void galaxies, it varies between 3 and 8. Similar studies conducted using the SDSS catalogue have set this limit at 20 (e.g., Ellison et al. 2008), but in our case, since the GAMA sample size is smaller, it does not allow for the creation of such large control samples.
@@
These criteria allow us to obtain a control sample of 2 114 field galaxies versus the 2 068 galaxies in filaments. The ECDFs for the stellar mass and colour of both the control sample and galaxies in filaments are shown in Figure 2. Furthermore, since all the p-values collected in Table 2 are greater than the significance level, = 0.05, it could be assumed that the control sample for galaxies in filaments and the sample of galaxies in filaments follow the same distribution.
@@
In addition, it is important to highlight that the KS and AD tests are designed for 1-dimensional samples, and their extension to higher dimensions can lead to issues. Thus, we tested our control samples by applying the 2-dimensional KS test using the mkstest package from Python (Naaman 2021). On the other hand, note that the upper limit of 0.25 for some of the p-values obtained from the AD test is due to the limitations of the Python package used (scipy package; Virtanen et al. 2020).
@@ 3 Results
Once the control samples have been determined for each sample of galaxies (filaments, tendrils, and voids), the – and –SFR relations can be studied. Specifically, the goal will be to establish if there are differences in metallicity, , or star formation rate, SFR, between the galaxy samples and their corresponding control samples. To quantify these differences, we will fit the relations using two methods: least-squares (Subsection 3.1) and Bayesian fitting (Subsection 3.2).
@@ 3.1 Least-squares method
Figure 5 presents the least-squares fits for the –SFR and – relations using the least-squared method. A linear regression was applied to the –SFR relation, while a third-order polynomial fit was used for the – relation. Additionally, the residuals of the fit approach zero, indicating a robust and well-constrained fit.
For the – relations, the differences in metallicity between the galaxies of each type and their corresponding control samples are small. However, for the –SFR relations, the SFR differences in SFR become appreciable (see Figure 5). To robustly analyse these results and draw conclusions, it is essential to consider the associated errors of these differences. Table 3 summarises the differences in and SFR found and their associated errors calculated using Bootstrap.
@@ 3.2 Bayesian method
Bayesian regression is a type of conditional modelling where a linear combination of other variables describes the mean of a variable. The goal is to obtain the posterior probability distributions of the regression coefficients. This method has been implemented here by creating models in Stan (Carpenter et al. 2017) using the R programming language. Error estimates are accounted for in these regression models. Therefore, making a Markov Chain Monte Carlo (MCMC) converge diagnostic makes it possible to obtain robust fits for the scaling relations.
Figure 5 shows the Bayesian fits obtained for the – and –SFR relations. Similar to the least-squares method, the – relation shows small differences in metallicity between the galaxies of each type and their corresponding control samples. In contrast, for the –SFR relations the SFR differences become significant (see Figure 5). Table 4 summarises the differences in metallicity and SFR found, along with their associated errors calculated using the standard deviation of the Markov Chains.
@@
It is interesting to see how, for the results summarized in Figure 3, the Bayesian results are slightly higher compared to the least-squares ones for the SFR, and the opposite for Z. These small differences are withing the errors, and both are consistent with each other, showing the robustness of our results using either method. Nevertheless, one of the methods could lead to an under- or overestimation. Even so, using the KS and AD tests to compare the and SFR distributions of the problem and control samples, we confirm again that the differences found are statistically significant, thus validating the differences found.
Finally, Figure 6 shows the Bayesian results for and SFR as full 2-dimensional Bayesian distributions. This approach avoids reducing the results to single values with associated errors and provides a 2-dimensional probability distribution instead. Note that there is no correlation between both and SFR differences.
@@ 3.3 Non-grouped galaxies
In this subsection, we analyse galaxies located in filaments or tendrils but not associated with any group, aiming to assess the direct impact of the environment on their properties. To achieve this, we use the GAMA group catalogue (G3C) Robotham et al. (2011) to identify filament and tendril galaxies that do not belong to any galaxy group.
With this information, we would be able to assess whether the evolution of these parameters is more dependent on large-scale structures or local-scale ones.
A control sample, which passed the KS and AD tests (see p-values in Table 2), was created for these non-grouped filament galaxies. A total of 1 435 field galaxies were found compared to 1 051 galaxies in the non-grouped filament galaxies. The same process was applied to non-grouped galaxies in tendrils. In this case, a total of 1 499 field galaxies were identified, compared to 1 143 non-grouped galaxies in tendrils.
@@ 3.4 Morphologies
Finally, in this section, we analyse the morphology of filaments, non-grouped filaments, tendrils, non-grouped tendrils, and void galaxies with their respective control samples. Our goal is to see if there are morphological differences that could be attributed to the environment. For this analysis, we use the GAMA Sèrsic catalogue as an indicator of morphology, which provides the Sèrsic index in multiple bands (Kelvin et al. 2012).
@@
The errors of the medians (shaded regions in Fig. 7) for each distribution overlap with each other. On the other hand, when applying the KS and AD tests to these distributions, they are only satisfied for the void galaxies compared to their control sample.
This implies that, although the medians of the distributions are similar, small morphological differences exist. For instance, the shape of the distribution of filaments and tendrils is slightly skewed towards early-type galaxies.
@@ 4 Discussion
This work analyses the – and –SFR relations of a selected sample of star-forming galaxies extracted from the GAMA survey (Driver et al. 2011; Liske et al. 2015). The main objective here is to identify differences in and SFR based on the type of galaxy according to the environment (filaments, tendrils, and voids), using generated control samples for each type. To achieve this, Bayesian and classical statistics were used. In this way, we investigate how different extreme environments, from voids to filaments, affect galaxy properties.
@@
Our analysis reveals variations in and SFR across different environments. In particular, galaxies residing in filaments exhibit a higher . While their history of undergoing multiple star-forming cycles may contribute to this enhancement (Ibrahim & Kobayashi 2024; Goswami et al. 2024), it is not clear how the nature of the filamentary environment itself plays a dominant role. It is known that dense environments facilitate the accretion of metal-enriched gas, boosting the metallicity of galaxies in these environments relative to those in less dense regions (Aracil et al. 2006; Cooper et al. 2008). Furthermore, these phenomena align with the SFR they are associated with, as the feedback environmental processes, combined with the effects of galactic interactions (Das et al. 2023), eject the gas that is favourable for incrementing significantly the SFR (Pandey et al. 2024). On the other hand, tendril galaxies exhibit a slightly more enhanced SFR and similar in comparison to galaxies in filaments. This suggests that both structures exhibit similar behaviour owing to their small differences. In contrast, void galaxies exhibit the lowest , which has already been observed in other studies, linking this phenomenon to gas accretion by the halo and their SFH (Domínguez-Gómez et al. 2023a). However, these galaxies show the highest SFR from all the studied samples. This can be explained by continuous star formation episodes (Beygu et al. 2016; Moorman et al. 2016; Domínguez-Gómez et al. 2023b), i.e., short recurrent star formation bursts arising from cold gas reserves from the halo. Furthermore, the low-density environments in which these objects are immersed prevent other galaxies from stripping their gas, thus allowing for ongoing star formation (Domínguez-Gómez et al. 2022; Argudo-Fernández et al. 2024).
With all this information, we observe that the and SFR differences obtained (Figure 3) for galaxies in filaments, tendril and void galaxies are consistent with the results of previous studies (Ellison et al. 2008, e.g.,; Sotillo-Ramos et al. 2021; Hoosain et al. 2024).
@@
In filaments and tendrils, interactions between galaxies can trigger star formation. In contrast, in voids, star formation is sustained due to the isolation of galaxies, and the surrounding halo may be accreted, contributing to the enhancement of the SFR. This increase in SFR as one moves toward lower dense environments aligns with the observed decrease in in the opposite density direction, which can be attributed to the accretion of enriched gas and interactions for galaxies in filaments and tendrils, and primordial gas accretion from the galaxy’s halo for void galaxies.
@@
On the other hand, galaxies in filaments and tendrils that do not belong to any galaxy group were studied in Subsection 3.3. For these non-grouped galaxies, only minimal differences in and SFR were found compared to their control sample of field galaxies (see Figure 3). These results suggest that the evolution of ungrouped galaxies, even if they are part of filaments or tendrils, is driven without important external perturbations, leading to a slower and more regulated gas consumption (Crone Odekon et al. 2018).
@@
The study of these ungrouped galaxies is particularly important in our work for two reasons. First, these galaxies exhibit minimal differences in and SFR compared to field galaxies, being largely compatible with them in most cases. Second, these differences are significantly smaller than those observed in the overall population of galaxies in filaments and tendrils, which includes both ungrouped and grouped galaxies. This suggests that the grouped galaxies are responsible for the increase in and SFR. Therefore, the values of and SFR are more influenced by the local density (e.g., group, cluster) in which the galaxy resides, rather than by the large-scale structure (filament, tendril, or void). Consequently, the evolutionary properties of a galaxy are determined primarily by the density of its local environment. This is in agreement with other studies, such as those by Kauffmann et al. (2004), Vulcani et al. (2012), Robotham et al. (2013), Hoosain et al. (2024), and O’Kane et al. (2024). Hence, large-scale structure can be understood as a second-order factor that significantly influences the characteristics of galaxies.
@@
In the context of the nature vs. nurture debate, our results suggest that ”nurture”, specifically local density, has a stronger impact on galaxy evolution than mere membership in filaments or tendrils. However, we note that ”nature,” or the intrinsic properties of galaxies such as internal feedback, gas accretion, disk instabilities, etc. (e.g., Tanaka et al. 2004; Maltby et al. 2010; Hughes et al. 2013; Silva-Lima et al. 2022), also plays a significant role, though this aspect is not explored in the present study.
@@
Finally, regarding the morphologies studied in section 3.4, Figure 7 shows that the errors of the medians for each morphological distribution overlap. When applying the KS and AD tests, they are only satisfied for the void galaxies in comparison to their control sample. This suggests that, although the medians of the distributions are nearly identical, there are small morphological differences in the direction of early-types. This has already been analysed in previous studies, such as that by Porter et al. (2023).
@@
Furthermore, a morphologically driven quenching can be ruled out for SF galaxies, as the Sérsic index distribution of galaxies in filaments, tendrils, and voids is similar to that of their corresponding control samples (Ricciardelli et al. 2017, e.g.,; Beygu et al. 2017; Smith et al. 2022). Lastly, it is worth noting that there does not exist any correlation between both and SFR differences (see Figure 6).

### v2

@@ 2.4 Galaxies in filaments, tendrils, and voids
For this work, we used the classification provided by Alpaslan et al. (2014a), who applied the FoF algorithm (Merson et al. 2013) to classify the galaxies in the GAMA survey. This classification divides galaxies into the following types.
Galaxies in filaments. This classification includes all galaxies within an orthogonal distance of 4.12 Mpc from filaments. Therefore, they are situated in dense environments and are susceptible to the influences of their surroundings.
Tendril galaxies. Included in this category are all galaxies within an orthogonal distance (4.56 Mpc) from tendrils, and simultaneously located beyond a minimum distance of 4.12,Mpc from any filament. Tendrils are slender, thread-like extensions that serve as connecting links between certain filaments. This term was first introduced by Alpaslan et al. (2014b).
Void galaxies. These are galaxies that are more than 4.56 Mpc from both filament and tendril structures. As a result, they are situated in largely underdense regions called voids, where they remain unaffected by external influences.
Considering these different types of galaxies, the sample of SF galaxies used in this study was classified into these three types, by matching the GAMA catalogue and the catalogue of filaments (Alpaslan et al. 2014a), as shown in Table 3.
Filaments | 2 068 | ||
Tendrils | 1 466 | ||
@@
From the initial 13 689 SF galaxies, 15% correspond to galaxies in filaments, 11% to tendril galaxies, and 1% to void galaxies. The percentage of the remaining unclassified galaxies corresponds mostly to field galaxies and will be used in the following section to create the control sample.
@@ 2.5 Control samples
We constructed such a control sample using field galaxies, selected to serve as a reference population. Field galaxies are identified from the original GAMA spectroscopic catalogue by excluding all galaxies in filaments, tendrils, and voids. In addition, we also removed galaxies that are members of groups or pairs. These exclusions are based on classifications from the GAMA filament catalogue (Alpaslan et al. 2014a), and the GAMA Galaxy Group Catalogue (G3C) (Robotham et al. 2011). We note that we worked only with SF galaxies with the same S/N and redshift cuts described previously in this section. By applying these criteria, we obtained a well-defined and uncontaminated sample of field galaxies for our analysis.
@@
Previous studies investigating environmental effects have generated control samples using similar methodologies and criteria (Ellison et al. 2008, e.g.; Scudder et al. 2012; Chen et al. 2020; Garduño et al. 2021; Sotillo-Ramos et al. 2021). We note that field galaxies are used as a control sample because they are not exposed to extreme environments, that is, they are neither located in structures such as filaments or tendrils, nor highly underdense regions such as voids. This makes them suitable for defining control samples, as done in previous studies (e.g. Ellison et al. 2008).
@@
To ensure the validity of our study, these field galaxies must share similar characteristics with those in filaments, tendrils, and voids. We followed previous works and used stellar mass and colour as our main criteria (Ellison et al. 2008, e.g., 2011, 2013), since these two parameters robustly characterise the main properties of a galaxy (Baldry et al. 2004; Peng et al. 2010). Given that our sample of galaxies is relatively local (z 0.3), the evolutionary effects are negligible (e.g. Lara-López et al. 2009). Therefore, we ensure that the differences in stellar mass and colour between galaxies in filaments, tendrils, and voids and those in the control sample remain minimal, adhering to specific selection criteria.
@@
These selection criteria, based on stellar mass and colour, were established using the Kolmogorov-Smirnov (KS) and Anderson-Darling (AD) tests as references. These tests assess the null hypothesis that two samples originate from the same population without requiring the specification of the distribution function of that population. The KS test is good enough to verify whether the created control sample and the sample of galaxies in filaments, tendrils, or voids can be assumed to be drawn from the same galaxy population. However, we complement this test with the AD test since it is more sensitive to the tails of the Empirical Cumulative Distribution Function (ECDF), thus ensuring that the samples are similar. The KS and AD tests accept the null hypothesis that both samples can be assumed to be drawn from the same population if the p-value is higher than the significance level = 0.05.
@@
During the creation of the control sample, additional considerations include limiting the number of matched field galaxies per target galaxy. For galaxies in filaments, a maximum of 20 field galaxies per filament galaxy was set. For tendril galaxies, this limit ranges from 2 to 20, while for void galaxies, it varies between 3 and 8. Similar studies conducted using the SDSS catalogue have set this limit at 20 (e.g. Ellison et al. 2008), but in our case, since the GAMA sample size is smaller, it does not allow for the creation of such large control samples.
@@
These criteria allow us to obtain a control sample of 2 114 field galaxies versus the 2 068 galaxies in filaments. The ECDFs for the stellar mass and colour of both the control sample and galaxies in filaments are shown in Figure 2. Furthermore, since all the p-values collected in Table 2 are greater than the significance level, = 0.05, it can be assumed that the control sample for galaxies in filaments and the sample of galaxies in filaments follow the same distribution.
@@
In addition, it is important to highlight that the KS and AD tests are designed for one-dimensional samples, and their extension to higher dimensions can lead to issues. Thus, we tested our control samples by applying the two-dimensional KS test using the mkstest package from Python (Naaman 2021). On the other hand, note that the upper limit of 0.25 for some of the p-values obtained from the AD test is due to the limitations of the Python package used (scipy package; Virtanen et al. 2020).
@@ 3 Results
Once the control samples have been determined for each sample of galaxies (filaments, tendrils, and voids), the – and –SFR relations can be studied. Specifically, the goal is to establish if there are differences in metallicity, , or star formation rate, SFR, between the galaxy samples and their corresponding control samples. To quantify these differences, we will fit the relations using two methods: least-squares fitting (Sect. 3.1) and Bayesian fitting (Sect. 3.2).
@@ 3.1 Least-squares method
Figure 5 presents fits for the – and –SFR relations using the least-squares method. A third-order polynomial fit was used for the – relation, while a linear regression was applied to the –SFR relation. Additionally, the residuals of the fit approach zero, indicating a robust and well-constrained fit.
For the – relations, the differences in metallicity between the galaxies of each type and their corresponding control samples are small. However, for the –SFR relations, the SFR differences in SFR become appreciable (see Figure 5). To robustly analyse these results and draw conclusions, it is essential to consider the associated errors of these differences. Table 3 summarises the differences in and SFR found and their associated errors calculated using bootstrapping.
@@ 3.2 Bayesian method
Bayesian regression is a type of conditional modelling where a linear combination of other variables describes the mean of a variable. The goal is to obtain the posterior probability distributions of the regression coefficients. This method was implemented here by creating models in Stan (Carpenter et al. 2017) using the R programming language. Error estimates are accounted for in these regression models. Therefore, using a Markov Chain Monte Carlo (MCMC) convergence diagnostic makes it possible to obtain robust fits for the scaling relations.
Figure 5 shows the Bayesian fits obtained for the – and –SFR relations. Similarly to the least-squares method, the – relation shows small differences in metallicity between the galaxies of each type and their corresponding control samples. In contrast, for the –SFR relations the SFR differences become significant (see Figure 5). Table 4 summarises the differences in metallicity and SFR found, along with their associated errors calculated using the standard deviation of the Markov chains.
@@
For the results summarised in Figure 3, it is interesting that the Bayesian results are slightly higher compared to the least-squares results for the SFR, and the opposite for Z. These small differences are within the errors, and both are consistent with each other, showing the robustness of our results using either method. Nevertheless, one of the methods could lead to an under- or overestimation. Even so, using the KS and AD tests to compare the and SFR distributions of the problem and control samples, we confirm again that the differences found are statistically significant, thus validating the differences found.
Finally, Figure 6 shows the Bayesian results for and SFR as full two-dimensional Bayesian distributions. This approach avoids reducing the results to single values with associated errors and provides a two-dimensional probability distribution instead. We note that there is no correlation between the and SFR differences.
@@ 3.3 Non-grouped galaxies
In this section, we analyse galaxies located in filaments and tendrils but not associated with any group, aiming to assess the direct impact of the environment on their properties. To achieve this, we used the GAMA group catalogue (G3C) (Robotham et al. 2011) to identify filament and tendril galaxies that do not belong to any galaxy group. With this information, we were able to assess whether the evolution of these parameters is more dependent on large-scale structures or local-scale ones.
A control sample, which passed the KS and AD tests (see p-values in Table 2), was created for these non-grouped galaxies in filaments. A total of 1 435 field galaxies were found compared to 1 051 non-grouped galaxies in filaments. The same process was applied to non-grouped galaxies in tendrils. In this case, a total of 1 499 field galaxies were identified, compared to 1 143 non-grouped galaxies in tendrils.
@@ 3.4 Morphologies
Finally, in this section, we analyse the morphology of galaxies in filaments (grouped and non-grouped), tendrils (grouped and non-grouped), and voids with their respective control samples. Our goal is to see if there are morphological differences that could be attributed to the environment. For this analysis, we used the GAMA Sèrsic catalogue as an indicator of morphology, which provides the Sèrsic index in multiple bands (Kelvin et al. 2012).
@@
The errors of the medians (shaded regions in Fig. 7) for each distribution overlap with each other. On the other hand, when applying the KS and AD tests to these distributions, they are only satisfied for the void galaxies compared to their control sample. This implies that, although the medians of the distributions are similar, small morphological differences exist. For instance, the shape of the distribution of filaments and tendrils is slightly skewed towards early-type galaxies.
@@ 4 Discussion
In this work, we describe our analyses of the – and –SFR relations of a selected sample of star-forming galaxies extracted from the GAMA survey (Driver et al. 2011; Liske et al. 2015). The main objective here was to identify differences in and SFR based on the type of galaxy according to the environment (filaments, tendrils, and voids), using generated control samples for each type. To achieve this, Bayesian and classical statistics were used. In this way, we investigated how different extreme environments, from voids to filaments, affect galaxy properties.
@@
Our analysis revealed variations in and SFR across different environments. In particular, galaxies residing in filaments exhibit a higher . While their history of undergoing multiple star-forming cycles may contribute to this enhancement (Ibrahim & Kobayashi 2024; Goswami et al. 2024), it is not clear how the nature of the filamentary environment itself plays a dominant role. It is known that dense environments facilitate the accretion of metal-enriched gas, boosting the metallicity of galaxies in these environments relative to those in less dense regions (Aracil et al. 2006; Cooper et al. 2008). Furthermore, these phenomena align with the SFR they are associated with, as the feedback environmental processes, combined with the effects of galactic interactions (Das et al. 2023), eject the gas that is favourable for incrementing significantly the SFR (Pandey et al. 2024). On the other hand, tendril galaxies exhibit a slightly more enhanced SFR and similar in comparison to galaxies in filaments. This suggests that both structures exhibit similar behaviour owing to their small differences. In contrast, void galaxies exhibit the lowest , which has already been observed in other studies, linking this phenomenon to gas accretion by the halo and their SFH (Dom´ınguez-Gómez et al. 2023a). However, these galaxies show the highest SFR from all the studied samples. This can be explained by continuous star formation episodes (Beygu et al. 2016; Moorman et al. 2016; Dom´ınguez-Gómez et al. 2023b), i.e., short recurrent star formation bursts arising from cold gas reserves from the halo. Furthermore, the low-density environments in which these objects are immersed prevent other galaxies from stripping their gas, thus allowing for ongoing star formation (Dom´ınguez-Gómez et al. 2022; Argudo-Fernández et al. 2024). With all this information, we observe that the and SFR differences obtained (see Figure 3) for galaxies in filaments, tendrils and voids are consistent with the results of previous studies (Ellison et al. 2008, e.g.; Sotillo-Ramos et al. 2021; Hoosain et al. 2024).
@@
Overall, in filaments and tendrils, interactions between galaxies can trigger star formation. In contrast, in voids, star formation is sustained due to the isolation of galaxies, and the surrounding halo may be accreted, contributing to the enhancement of the SFR. This increase in SFR as one moves towards lower-density environments aligns with the observed decrease in in the opposite density direction, which can be attributed to the accretion of enriched gas and interactions for galaxies in filaments and tendrils, and primordial gas accretion from the galaxy’s halo for void galaxies.
@@
On the other hand, galaxies in filaments and tendrils that do not belong to any galaxy group were studied in Section 3.3. For these non-grouped galaxies, only minimal differences in and SFR were found compared to their control sample of field galaxies (see Figure 3). These results suggest that the evolution of ungrouped galaxies, even if they are part of filaments or tendrils, is driven without important external perturbations, leading to a slower and more regulated gas consumption (Crone Odekon et al. 2018).
@@
The study of these ungrouped galaxies is particularly important in our work for two reasons. First, these galaxies exhibit minimal differences in and SFR compared to field galaxies, being largely compatible with them in most cases. Second, these differences are significantly smaller than those observed in the overall population of galaxies in filaments and tendrils, which includes both ungrouped and grouped galaxies. This suggests that the grouped galaxies are responsible for the increase in and SFR. Therefore, the values of and SFR are more influenced by the local density (e.g. group, cluster) in which the galaxy resides, rather than by the large-scale structure (filament, tendril, or void). Consequently, the evolutionary properties of a galaxy are determined primarily by the density of its local environment. This is in agreement with other studies, such as those by Kauffmann et al. (2004), Vulcani et al. (2012), Robotham et al. (2013), Hoosain et al. (2024), and O’Kane et al. (2024). Hence, large-scale structure can be understood as a second-order factor that significantly influences the characteristics of galaxies.
@@
In the context of the ‘nature’ versus ‘nurture’ debate, our results suggest that nurture, specifically local density, has a stronger impact on galaxy evolution than mere membership in filaments or tendrils. However, we note that nature, or the intrinsic properties of galaxies, such as internal feedback, gas accretion, or disk instabilities also plays a significant role (Tanaka et al. 2004, e.g.; Maltby et al. 2010; Hughes et al. 2013; Silva-Lima et al. 2022), though this aspect is not explored in the present study.
@@
Finally, regarding the morphologies studied in Section 3.4, Figure 7 shows that the errors of the medians for each morphological distribution overlap. When applying the KS and AD tests, they are only satisfied for the void galaxies in comparison to their control sample. Although the medians of the distributions are nearly identical, this suggest that there are small morphological differences in the direction of early types. This has already been analysed in previous studies, such as that by Porter et al. (2023).
@@
Furthermore, a morphologically driven quenching can be ruled out for SF galaxies, as the Sérsic index distribution of galaxies in filaments, tendrils, and voids is similar to that of their corresponding control samples (Ricciardelli et al. 2017, e.g.; Beygu et al. 2017; Smith et al. 2022). Lastly, it is worth noting that there is no correlation between the and SFR differences (see Figure 6).